### v1

# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

Abstract
Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained. Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets. We prompt GPT-4 to suggest expressions from data, which are then optimized and evaluated using external Python tools. These results are fed back to GPT-4, which proposes improved expressions while optimizing for complexity and loss. Using chain-of-thought prompting, we instruct GPT-4 to analyze the data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions. We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation. GPT-4 successfully rediscovered all five equations, and in general, performed better when prompted to use a scratchpad and consider scientific context. We also demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data. Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
Keywords Symbolic Regression  $\cdot$ Large Lanugae Models  $\cdot$ GPT-4
1 Introduction
Data analysis is ubiquitous in all disciplines, where identifying correlations between variables is key to finding insights, informing conclusions, supporting hypotheses, or developing a new theory. For scientific data, we often aim to find expressions with few adjustable parameters explaining the data while ensuring that they align with theory. Symbolic regression is a machine learning technique that approaches equation-based scientific discovery – given a dataset, it searches through some “space of possible equations” and identifies those that balance accuracy and simplicity. It is different from conventional regression methods, as symbolic regression infers the model structure from data rather than having a predetermined model structure.
Mathematically, symbolic regression is formulated as some form of optimization, not just of the constants in an equation, but as a search through “equation space" for optimal expressions. In this way, symbolic regression is a form of machine learning – as data is received, an internal model is updated to match the data; when the model fits the data well and can make predictions about unseen data, the algorithm is said to have “learned" the underlying patterns in the data. In contrast to popular machine learning algorithms like neural networks, symbolic regression does not just only fit the constants in an equation but also finds functional forms that match the data.
Symbolic regression methods mainly use genetic algorithms [1, 2, 3, 4] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached. Other approaches include using Markov chain Monte Carlo (MCMC) sampling [5, 6], mixed integer nonlinear programming [7, 8, 9], greedy tree searches [10], pre-trained transformer-based models [11, 12], and sparse matrix algorithms [13, 14, 15]. These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning. Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29]. As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery. LLMs are machine learning models adept at understanding and generating natural language. At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [30]. The term “large" in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained. Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks. However, GPT-3 demonstrated that language models can excel at tasks using in-context learning without necessitating fine-tuning [31, 32]. LLMs are now commonly used for tasks like chat, code generation, summarization, translation, etc. -- and quite remarkably, these tasks can be effectively accomplished by using English language as model input without the need for machine learning expertise. We have firsthand experience with these models’ capabilities by interacting with AI chatbots on platforms like ChatGPT, Claude, or Gemini. LLMs have a wide understanding of the world from their training data and can even solve simple math problems expressed in natural language [33, 34, 35]. They are contributing significantly in education and research [36, 37, 38, 39] medicine [40, 41], physical and social sciences [42, 43, 44], as well as in legal [45, 46], business [47, 48], and entertainment [49, 50, 51, 52] sectors.
Existing transformer-based symbolic regression programs [53, 11, 12] use models pretrained on large databases of synthetically-generated dataset/expression pairs, designed specifically for symbolic regression tasks. Thus, these approaches learn to pattern-match between datasets and math expressions, but they don’t employ iterations like those in genetic programming that optimize expressions for complexity and loss. In that respect, LLMs have been used to imitate evolutionary algorithms (EA) [54, 55, 56, 57, 58, 59, 60], and have specifically been applied to solve symbolic regression (SR) problems [56, 59]. Meyerson and coworkers [56] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators. The second approach more closely resembles our work, but what we do is even simpler: we task GPT-4 to generate and/or transform expressions freely. Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [56, 59], the context employed is limited to a list of previously-obtained expressions. We propose to expand the context to include data, as well natural language descriptions of the scientific context of the problem. To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [61, 62, 63]. Therefore, we also incorporate zero-shot chain-of-thought prompting with a scratchpad [64] to frame equation generation for symbolic regression as a reasoning problem [65] in the context of the data and free-form scientific information.
Symbolic regression requires equation generation and precise fitting of numerical constants. Yang and coworkers [66] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without showing the LLM the analytical form. They find that “LLM can often calculate the solution directly from the analytic form," however we find LLMs to be unreliable and inefficient for such tasks. Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [58], iteratively refining prompts for more accurate output. In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants.
This work aligns closely with two recent works [67, 68] in the literature in terms of methodology in that LLMs were used to generate expressions, and iterations were performed separately for optimization. Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using data and context in natural language as model input, and implementing a scratchpad to record model output, which can help reveal if there was a case of “test set leakage.” Merler and coworkers [67] used multimodal LLMs with image and text inputs (without context and/or scratchpad), but their benchmark datasets does not reflect scientific data. Conversely, Shojaee and coworkers [68] specifically focused on scientific problems (without data and/or scratchpad), and while their expression generation utilized cleaner Python functions it is difficult to evaluate if GPT-4 has encountered these questions as they do not include data but present the task as a problem similar to college homework assignments.
2 Methods
Preliminary tests
In our initial tests, we asked GPT-4 to generate mathematical expressions from scientific data. We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1). In all cases, the models produced expressions while “hallucinating” arbitrary coefficients. We revised the prompts to ask the LLMs to “show all steps" to gain insight into how the model selects these values [63]. In response, the output either provided Python code for optimization or, at times, mathematical steps for optimization.
Figure 1: Illustration of GPT-4 attempting symbolic regression. GPT-4 predicts expressions with optimized coefficients when passed a dataset for nitrogen adsorption on mica [69]. The Python code snippet from Prompt 2 output has been truncated to keep the figure concise. Note that the actual parameter values produced by running the code differ from what GPT-4 generates. Figure 10 in SI shows the output from GPT-3.5 turbo which produces similar results.
Although the generated code accurately performed curve-fitting, the LLM hallucinated incorrect coefficients. Nonetheless, the generated remarks about the data patterns hinted that they may have the potential to generate accurate functional forms that can be optimized outside the LLM.
System design
Therefore, we designed a workflow (Figure 2) where we task GPT-4 with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM. A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE). These results are stored as a dictionary, the text of which is passed back to GPT-4 in a subsequent prompt asking to suggest better expressions. We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4, more frequently generating expressions that didn’t parse.
Figure 2: Workflow for using LLMs for SR. First, the dataset is sent to GPT-4 (with or without context) which is instructed to suggest expressions without optimizing parameters. The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [70] method with basin-hopping as the numerical optimizer [71]). Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions. The top performing expressions are sent as feedback to GPT-4, which is asked to suggest better expressions optimizing for both complexity and loss. The feedback loop is run for a set number of iterations.
We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format. We sort the expressions in descending order of mean squared error, then share this list as feedback, asking GPT-4 to suggest new equations optimizing for complexity and loss. In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation. We do not use chat history or any advanced forms of memory [72, 73, 74, 75] in this workflow. The history of generated expressions is maintained externally and provided as feedback in the iteration prompt. A Python function maintains this feedback loop by sorting and filtering a list of dictionaries based on MSE and complexity. At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window.
By making each call to GPT-4 independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
We can get the most out of an LLM by providing strategic text in the prompt. Prompt engineering is an art that involves structuring prompts that guide the model toward generating desired outputs. However, there is no one-size-fits-all method for crafting optimal prompts, as outcomes depend on the specific task and model. Even minor changes in wording or structure can influence the model’s output. Various prompting guidelines have been explored in the literature [78, 79, 80, 81], including providing clear instructions, emphasizing relevant context using text delimiters, breaking tasks into multiple steps, and incorporating problem-solving conditions. We implemented strategies like incorporating examples and a scratch pad and subsequently refined the prompts.
Prompt Engineering
We prototyped our system by testing its ability to rediscover the Langmuir adsorption isotherm ($q=c_{1}*p/(c_{2}+p)$) from experimental data [69]. This enabled us to quickly identify major structural improvements to the workflow; we further tailored the prompts while testing on more difficult problems.
Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from GPT-4 to ensure the SciPy function runs without any errors. A simple way to illustrate the expected outcome was by providing examples in a few-shot context [31, 82]. While this led to expressions matching the required syntax, we noticed the generated expressions resembled the examples we provided (Figure 3). While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax. The revised prompt (Prompt 2) eliminates this bias from the equation prediction process.
This motivated our two-prompt setup, with an initial prompt tasking GPT-4 to generate unbiased expressions in LaTeX (which it generated more reliably than as Python strings), and an iteration prompt receiving previously-generated examples as Python script (now without bias). To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating. Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model.
Recording analysis in a scratchpad: Studies have shown that LLM performance can be improved by slowing down the model or breaking down its tasks into smaller steps [61, 62]. One popular strategy is the “scratch pad" technique [64], which mimics how we solve problems by jotting down notes before presenting a final answer in exams. We implemented this in our workflow, instructing GPT-4 to generate responses in two parts: data analysis and observations in a scratch pad, followed by its conclusions. After implementing this technique, the model immediately generated higher-quality expressions (Figure 4).
(a) Effect of scratchpad on Langmuir Dataset
(b) Effect of scratchpad on Kepler Law Dataset
Figure 4: Illustration of scratchpad approach. We observe substantial, qualitative improvements in the predicted expressions after implementing the scratchpad technique (Prompt 2). The suggested expressions for (a) Langmuir’s and (b) Kepler’s Law dataset include operators (/// and sqrt, respectively) present in the target models ($y=\frac{c_{1}*x}{c_{2}+x}$ and $y^{2}=c_{1}x^{\frac{3}{2}}$, respectively)
Preventing redundant expressions: GPT-4 often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different. However, since the constants are yet to be fitted, these expressions are the same from a symbolic regression perspective. While a computer algebra system like SymPy [83] could in principle catch some redundant expressions by simplification to a canonical form [6, 28], this wouldn’t distinguish “SR-similar" expressions that become equivalent after fitting constants. Instead, we used prompt engineering to guide generation toward unique expressions: we added a note in the iteration prompt with examples showing how expressions in symbolic regression are similar before parameters are optimized. While this didn’t completely resolve the issue, we did observe a reduction in occurrences, and at times, the scratchpad revealed GPT-4 correctly addressing this by taking these examples into account (see Figure 5).
Figure 5: Illustration of GPT-4 outputs for SR-similar expressions.
Avoiding uninteresting expressions: During the iterative runs, GPT-4 attempted to improve its accuracy by repeatedly adding linear terms to suggested expressions from previous iterations. To address this issue, we encouraged the model to explore diverse expressions in the prompt. Additionally, in cases involving datasets with multiple independent variables, GPT-4 sometimes recommended excluding variables that exhibited weak correlation with the overall dataset pattern. While this may be useful in some contexts, we wanted expressions that made use of all of the available data, so we explicitly instructed the use of all variables. Additional constraints we implemented included limiting the types of math operators to include and preventing generation of implicit functions, as shown in Figure 6.
Consider scientific context: Our primary motivation for building this system was to test whether providing scientific context could shape the expressions generated by the LLM. SR programs often successfully generate expressions that fit the data well and are simple, yet they may not adhere to scientific principles or be otherwise “meaningful.” Yet scientists often have valuable insights into their domain that extend beyond these constraints, and they may not always know which specific expressions will best capture the nuances of their dataset or if entirely new expressions might be more effective. By incorporating scientific context, we aim to align the equation search to be consistent with scientific theories.
Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g. by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [84, 85, 86, 24, 18, 25, 87]. Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 11 in SI lists the context provided for all the datasets.
Figure 6: Illustration of GPT-4 following restrictions from prompt.
3 Results
We evaluate our workflow using experimental datasets associated with meaningful scientific context. SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data. From here, we selected three experimental datasets from astronomical observations (Bode’s Law, Hubble’s Law, and Kepler’s Law). To these, we added two experimental chemistry datasets: Langmuir and Dual-site Langmuir adsorption isotherm models [9].
Figure 12 in SI illustrates all the datasets with the target expression for each. We also tested our workflow using a dataset on friction losses in pipe flow [89]. This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6]. We included the whole dataset for each problem in the prompt to GPT-4 at the start and in each iteration. Because the Hubble and Leavitt’s datasets had an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to GPT-4 but used the original dataset when optimizing constants in SciPy. This reduced the tokens (and cost) of running GPT-4. We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
We conducted eight sets of experiments on each of the six datasets, running each 5 times to evaluate robustness. In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data. Empirical relations often involve field-specific operators, so the equation search should account for this. We refer to this set of tests as an “easy search" because the space of possible equations is constrained to that generated by the operators in our dataset. In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search.
Temperature is a hyperparameter used in stochastic models like LLMs to regulate the randomness of the model output [90, 91]. It adjusts the probabilities of the predicted words in the softmax output layer of the model. Lowering temperature favors words with higher probability, so when the model randomly samples the next word from the probability distribution, it will be more likely to choose a more predictable response. We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the datasets.
Figure 7: GPT-4 results on “easy" and “hard" searches. Easy search includes only basic operators while harder one adds more. All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on. The score at each iteration is a total of x/5 independent runs where target model expressions were found. With these settings; dual-site Langmuir was never found after 50 iterations; with a modified feedback loop, it was only found 1/5 times in easy searches with and without tools.
Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which is often a first guess when few operators are made available. However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or the context inhibited performance. This dataset is particularly noisy, as well; including the noisy data as context actually inhibits the search.
Bode: The results for Bode’s Law indicate that, similar to other SR programs [88], GPT-4 encounters difficulty accurately rediscovering the target model expression. Curiously, it performs best when the context is excluded and worst when the data is excluded. This suggests that the context we provided was counterproductive for the search, and purely reasoning about the data led greater success. We also note that GPT-4 finds the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$ when all operators are provided, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
Kepler: We find strong evidence that the prompt has triggered GPT-4’s memorization of Kepler’s Law: the scratchpad reveals GPT-4 associates the variable names in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification. Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in GPT-4’s training data than the other relationships. The hard search for Kepler’s Law with all tools on has led GPT-4 to explore more complex expressions since the target model complexity is only 5 with MSE of 46.6886. GPT-4 also finds the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$. However, SciPy optimizes and fits $c_{2}$ with to a floating point power ($\approx 1.5$). We did not consider this to be a rediscovery of the target model as such expressions are not dimensionally consistent.
Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round. In easier searches, GPT-4 consistently found it within 15 iterations; inclusion of context and data seemed to improve performance. However, the hard search was much more challenging, with Langmuir only discovered once, in the case with all tools on.
Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions. In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 8). However extrapolation shows this model is not theoretically correct (Figure 13 in SI).
This dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92]. Because this expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search. We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop). We found that when using basic operators and context, GPT-4 was getting close to finding the target model (Figure 8) upon running 15 iterations with this feedback loop. To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed. Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR. But when targeting models like dual-site Langmuir, these instructions may have been a liability, and we needed to adjust the prompt and feedback sent to GPT-4 to allow for the exploration of more complex expressions.
Figure 8: Pareto fronts for dual-site Langmuir dataset. The black line represents the best total front from the five independent runs. The target model is labeled as a blue star.
Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to GPT-4 exceeded the token limit. Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations. So, we developed a cost-saving scheme: only send a portion of the data in the prompt to GPT-4, while fitting and evaluating the generated expressions using the whole dataset (Figure 14 in SI).
Since GPT-4 generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging. We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions. So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error. Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to GPT-4 for feedback. To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6]. We modified the basic prompt to encourage GPT-4 to explore longer expressions and also tested the effect of “prodding" GPT-4 by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better. We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) random data points and the second (S2) with another random 36 (20%) data points. Overall, feeding in more data generated better-fitted and more complex expressions. Figure 9 shows the top expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^).
Figure 9: Models for Nikuradse dataset from GPT-4. Here P refers to prompt versions, and S refers to dataset. We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig. 9 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways. However it generalizes better with fewer data for prompt version 3 where we challenge GPT-4 by providing information about the MAE and complexity of BMS model.
We found an optimal expression from GPT-4 with a complexity of 41 and an MAE of 0.01086 (Figure 9 C). The MAE is approximately three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS). BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392. However unlike GPT-4, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps to identify this expression. Our GPT-4 model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions. We additionally assessed BMS on its default move probabilities and using 40 temperatures for parallel tempering as mentioned in the paper, and ran it for 153 MC steps with Nikuradse data -- it generated a constant function, and upon running it longer (1000 MC steps), it suggests an expression with MAE 0.13436 and complexity 25. Thus, we can see that GPT-4’s incremental suggestions for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS. However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost. Using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
BMS samples expressions from a probability distribution. Therefore, running for a long time is expected to be characterized by equilibration, in which the expressions converge to a region in the stationary distribution of expressions, after which no significant improvement would be observed apart from continuing to explore the region of most likely expressions. Genetic algorithms like Eureqa [3] do not converge to a stationary distribution and nor does our approach. On the contrary, running a chatbot for a very long time leads to degradation of its answers as the context gets longer [76, 77], so, in principle, our method may exhibit similar degradation with very long runs. However, we expect this effect to be mitigated because we discard the majority of the context after each iteration, only passing the top and recent expressions.
We also show a model from Eureqa in [6] which was run with default operator penalties and selected for the best expression from at least $10^{13}$ ones. The result from the EFS [85] model (best fit selected from 100 runs) in [6] is comparable to our model with an MAE of 0.00941. EFS is based on sparse regression that uses a genetic algorithm to generate basis functions automatically. EFS is known to be highly efficient, delivering expressions in seconds.
4 Discussion and Conclusion
SR programs that optimize for speed aim to generate expressions quickly. In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively. The clearest cases of “leveraging the context” occurred when GPT-4’s first guess included the target model among three expressions. But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions. However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context. Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens.
In general, we found natural language to be a rather clumsy interface for controlling expression length. Different prompts and feedback mechanisms led to distributions of expressions with varying length. Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [93]. We ran separate, focused tests to evaluate the effect of prompting on expression length, and GPT-4 did not obey instructions that requested, for example, "expressions of length 17." Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length. The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting complexity.
Testing our approach on equation rediscovery using GPT-4 invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources. Indeed, the data are publicly available, as well, though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets. We found strong evidence of this because of our scratchpad implementation, which revealed when it was thinking of Kepler’s law before seeing the data. Nonetheless, we found evaluating the model outputs to be informative.
Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include. This context may include experimental details, instrument specifications, and literature, and even a “memorized” explanation of the data by an LLM or retrieved from context in a retrieval-augmented generation scheme [94] would be valuable. A true blind test would be to rediscover a novel scientific law using an LLM with a knowledge cutoff predating the seminal publication.
There are two ways to implement in-context learning in LLMs through prompts. One is the Few-Shot prompting method, where we condition the model on a few specific examples related to the task that helps the model understand and perform the task more accurately. The other is the Zero-Shot prompting method, where the output relies solely on template-based prompts without specific task examples, allowing the model to infer how to handle the task from general instructions. Our approach does not provide examples of symbolic regression procedures within the prompts. Instead, we guide the model to engage in freeform chain-of-thought reasoning about the context as it prepares suggestions for new equations.
Biasing the search space in a standard SR program can be challenging and requires advanced software and coding skills. Interdisciplinary work demands significant time and resources from researchers. Natural language interfaces in LLMs can help reduce some of these barriers by making program execution more accessible, even without expertise in software development. Well-crafted prompts empower language models to perform diverse tasks, allowing them to adapt to different contexts and objectives. We can guide these models through prompts to generate responses that align with our specific needs, whether solving complex problems, developing creative content, or analyzing data.
Acknowledgements
We thank Roger Guimerà for sharing the detailed results of all models on Nikuradse dataset. This material is based upon work supported by the National Science Foundation under Grant No. #2138938.
Conflicts of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The code, prompts, and data supporting the findings of this work are available at https://github.com/ATOMSLab/pySR_adsorption.
5 Supporting Information
Figure 10: Illustration of GPT-3.5 attempting to perform SR.
Figure 11: Context provided to GPT-4 for all the datasets
Figure 12: Datasets explored using GPT-4 for SR
Figure 13: GPT-4-generated model does not follow theory on extrapolation. It gives a negative value for adsorption loading at lower pressures.
Figure 14: Nikuradse Dataset. The red and yellow points represent the data sent to GPT-4 (selected randomly from the original dataset), while the grey ones show the original dataset sent to SciPy for optimization.

### v2

@@
 Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets.
-We prompt GPT-4 to suggest expressions from data, which are then optimized and evaluated using external Python tools.
-These results are fed back to GPT-4, which proposes improved expressions while optimizing for complexity and loss.
-Using chain-of-thought prompting, we instruct GPT-4 to analyze the data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions.
+We prompt GPT-4 and GPT-4o models to suggest expressions from data, which are then optimized and evaluated using external Python tools.
+These results are fed back to the models, which proposes improved expressions while optimizing for complexity and loss.
+Using chain-of-thought prompting, we instruct the models to analyze data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions.
 We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation.
-GPT-4 successfully rediscovered all five equations, and in general, performed better when prompted to use a scratchpad and consider scientific context.
-We also demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data.
+GPT-4 and GPT-4o successfully rediscovered all five equations and generally performed better when prompted to use a scratchpad and consider scientific context.
+GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse datasets.
+We demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data.
+We also observe how theory can sometimes offset noisy data and, in other cases, data can make up for poor context.
 Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
-Keywords Symbolic Regression  $\cdot$ Large Lanugae Models  $\cdot$ GPT-4
+Keywords scientific discovery  $\cdot$ GPT-4  $\cdot$ adsorption  $\cdot$ Nikuradse  $\cdot$ symbolic regression
 1 Introduction
@@
 These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning.
-Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29].
+Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
 As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
@@
 At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [39].
-The term “large" in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained.
+The term “large” in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained.
 Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks.
@@
 Meyerson and coworkers [31] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators.
-The second approach more closely resembles our work, but what we do is even simpler: we task GPT-4 to generate and/or transform expressions freely.
+The second approach more closely resembles our work, but what we do is even simpler: we task LLMs to generate and/or transform expressions freely.
 Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [31, 34], the context employed is limited to a list of previously-obtained expressions.
@@
 In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants.
-This work aligns closely with two recent works [67, 68] in the literature in terms of methodology in that LLMs were used to generate expressions, and iterations were performed separately for optimization.
-Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using data and context in natural language as model input, and implementing a scratchpad to record model output, which can help reveal if there was a case of “test set leakage.” Merler and coworkers [67] used multimodal LLMs with image and text inputs (without context and/or scratchpad), but their benchmark datasets does not reflect scientific data.
-Conversely, Shojaee and coworkers [68] specifically focused on scientific problems (without data and/or scratchpad), and while their expression generation utilized cleaner Python functions it is difficult to evaluate if GPT-4 has encountered these questions as they do not include data but present the task as a problem similar to college homework assignments.
+This approach aligns closely with two recent works [36, 37] in the literature where LLMs were used to generate expressions and optimizations were performed via iterations.
+Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using both data and context in natural language as model input, and implementing a scratchpad to record raw model output as it solves the problem.
+This promotes informal reasoning about the data and context, and helps reveal “test set leakage.”
 2 Methods
 Preliminary tests
-In our initial tests, we asked GPT-4 to generate mathematical expressions from scientific data.
+In our initial tests, we asked LLMs to generate mathematical expressions from scientific data.
 We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1).
@@
 System design
-Therefore, we designed a workflow (Figure 2) where we task GPT-4 with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM.
+Therefore, we designed a workflow (Figure 2) where we task GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM.
 A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE).
-These results are stored as a dictionary, the text of which is passed back to GPT-4 in a subsequent prompt asking to suggest better expressions.
-We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4, more frequently generating expressions that didn’t parse.
+These results are stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions.
+We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4 and GPT-4o, more frequently generating expressions that didn’t parse.
 Figure 2: Workflow for using LLMs for SR.
-First, the dataset is sent to GPT-4 (with or without context) which is instructed to suggest expressions without optimizing parameters.
+First, the dataset is sent to LLMs (with or without context) which is instructed to suggest expressions without optimizing parameters.
 The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [70] method with basin-hopping as the numerical optimizer [71]).
 Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions.
-The top performing expressions are sent as feedback to GPT-4, which is asked to suggest better expressions optimizing for both complexity and loss.
+The top performing expressions are sent as feedback to LLMs, which is asked to suggest better expressions optimizing for both complexity and loss.
 The feedback loop is run for a set number of iterations.
-We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format.
-We sort the expressions in descending order of mean squared error, then share this list as feedback, asking GPT-4 to suggest new equations optimizing for complexity and loss.
+We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 and GPT-4o to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format.
+We sort the expressions in descending order of mean squared error, then share this list as feedback, asking the LLMs to suggest new equations optimizing for complexity and loss.
 In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation.
@@
 At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window.
-By making each call to GPT-4 independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
+By making each call to LLMs independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
 We can get the most out of an LLM by providing strategic text in the prompt.
@@
 This enabled us to quickly identify major structural improvements to the workflow; we further tailored the prompts while testing on more difficult problems.
-Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from GPT-4 to ensure the SciPy function runs without any errors.
+Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors.
 A simple way to illustrate the expected outcome was by providing examples in a few-shot context [40, 82].
@@
 Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax.
-The revised prompt (Prompt 2) eliminates this bias from the equation prediction process.
-This motivated our two-prompt setup, with an initial prompt tasking GPT-4 to generate unbiased expressions in LaTeX (which it generated more reliably than as Python strings), and an iteration prompt receiving previously-generated examples as Python script (now without bias).
+The revised prompts (Prompt 2 and 3) eliminate this bias from the equation prediction process.
+This motivated our two-prompt setup, with an initial prompt tasking LLMs to generate unbiased expressions in LaTeX and an iteration prompt receiving previously-generated examples as Python script (now without bias).
 To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating.
 Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model.
+While running the GPT-4o model, we realized that most syntax errors stemmed from incorrect LaTeX formatting.
+To address this, we made a minor adjustment by removing the term "LaTeX" and allowing the model to generate expressions as Python strings instead.
+This small change was was sufficient to remove bias and was also less prone to syntax errors.
+We nervetheless processed the generated Python expressions through our formatting function, and it worked seamlessly.
 Recording analysis in a scratchpad: Studies have shown that LLM performance can be improved by slowing down the model or breaking down its tasks into smaller steps [63, 64].
 One popular strategy is the “scratchpad" technique [66], which mimics how we solve problems by jotting down notes before presenting a final answer in exams.
-We implemented this in our workflow, instructing GPT-4 to generate responses in two parts: data analysis and observations in a scratch pad, followed by its conclusions.
+We implemented this in our workflow, instructing LLMs to generate responses in two parts: data analysis and observations in a scratchpad, followed by its conclusions.
 After implementing this technique, the model immediately generated higher-quality expressions (Figure 4).
@@
 The suggested expressions for (a) Langmuir’s and (b) Kepler’s Law dataset include operators (/// and sqrt, respectively) present in the target models ($y=\frac{c_{1}*x}{c_{2}+x}$ and $y^{2}=c_{1}x^{\frac{3}{2}}$, respectively)
-Preventing redundant expressions: GPT-4 often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different.
+Preventing redundant expressions: The LLMs often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different.
 However, since the constants are yet to be fitted, these expressions are the same from a symbolic regression perspective.
@@
 Instead, we used prompt engineering to guide generation toward unique expressions: we added a note in the iteration prompt with examples showing how expressions in symbolic regression are similar before parameters are optimized.
-While this didn’t completely resolve the issue, we did observe a reduction in occurrences, and at times, the scratchpad revealed GPT-4 correctly addressing this by taking these examples into account (see Figure 5).
+While this didn’t completely resolve the issue, we did observe a reduction in occurrences, and at times, the scratchpad revealed the models would correctly address this by taking the examples into account (see Figure 5).
 Figure 5: Illustration of GPT-4 outputs for SR-similar expressions.
-Avoiding uninteresting expressions: During the iterative runs, GPT-4 attempted to improve its accuracy by repeatedly adding linear terms to suggested expressions from previous iterations.
+Avoiding uninteresting expressions: During the iterative runs, LLMs attempted to improve its accuracy by repeatedly adding linear terms to suggested expressions from previous iterations.
 To address this issue, we encouraged the model to explore diverse expressions in the prompt.
-Additionally, in cases involving datasets with multiple independent variables, GPT-4 sometimes recommended excluding variables that exhibited weak correlation with the overall dataset pattern.
+Additionally, in cases involving datasets with multiple independent variables, LLMs would sometimes recommend excluding variables that exhibited weak correlation with the overall dataset pattern.
 While this may be useful in some contexts, we wanted expressions that made use of all of the available data, so we explicitly instructed the use of all variables.
@@
 by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [84, 85, 86, 24, 18, 25, 87].
-Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 11 in SI lists the context provided for all the datasets.
+Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 17 in SI lists the context provided for each of the datasets.
 Figure 6: Illustration of GPT-4 following restrictions from prompt.
 3 Results
-We evaluate our workflow using experimental datasets associated with meaningful scientific context.
+We evaluated our workflow using experimental datasets associated with meaningful scientific context.
 SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data.
@@
 This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6].
-We included the whole dataset for each problem in the prompt to GPT-4 at the start and in each iteration.
-Because the Hubble and Leavitt’s datasets had an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to GPT-4 but used the original dataset when optimizing constants in SciPy.
-This reduced the tokens (and cost) of running GPT-4.
+We included the whole dataset for each problem in the prompt to LLMs at the start and in each iteration.
+Because the Hubble’s dataset has an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to LLMs but used the original dataset when optimizing constants in SciPy.
+This reduced the tokens (and cost) of running the models.
 We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
-We conducted eight sets of experiments on each of the six datasets, running each 5 times to evaluate robustness.
+We conducted eight sets of experiments for each model (GPT-4 and GPT-4o) on six datasets, running each set five times to evaluate its robustness.
 In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data.
@@
 In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search.
+All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on.
+In the following text, we discuss each dataset individually and analyze their performance.
+The score at each iteration is a total of x/5 independent runs where target model expressions were found.
 Temperature is a hyperparameter used in stochastic models like LLMs to regulate the randomness of the model output [90, 91].
@@
 We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the datasets.
-Figure 7: GPT-4 results on “easy" and “hard" searches.
-Easy search includes only basic operators while harder one adds more.
-All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on.
-The score at each iteration is a total of x/5 independent runs where target model expressions were found.
-With these settings; dual-site Langmuir was never found after 50 iterations; with a modified feedback loop, it was only found 1/5 times in easy searches with and without tools.
-Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which is often a first guess when few operators are made available.
-However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or the context inhibited performance.
-This dataset is particularly noisy, as well; including the noisy data as context actually inhibits the search.
-Bode: The results for Bode’s Law indicate that, similar to other SR programs [88], GPT-4 encounters difficulty accurately rediscovering the target model expression.
+Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which was often a first guess in the GPT-4 model when few operators are made available.
+However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or context inhibited performance (Figure 7).
+This dataset is particularly noisy, and including the noisy data as context actually inhibits the search as observed by consistent scores of 5/5 from both models when they were run without the data.
+Furthermore, the scratchpad in GPT-4o reveals that when the model first guesses Hubble’s Law, it then tends to reason and deduce a more complex, non-linear model in context of the noisy data.
+Figure 7: GPT-4 and GPT-4o results on Hubble’s data.
+In the GPT-4 model, all the trends except ‘Without Context’ overlap at score 5 for the “Easy Search” plot, and the trends ‘With All Tools’ and ‘Without Data’ overlap at score 4 for the “Hard Search”.
+In the GPT-4o model, the trends for ‘Without Data’ and ‘Without Context’ overlap at score 0 for the “Easy Search” plot, and all the trends except ‘Without Data’ overlap at score 0 for the “Hard Search”.
+Transparency and different marker styles are used to differentiate these trends.
+Bode: The results for Bode’s Law indicate that similar to other SR programs [88], LLMs encounter difficulty in accurately rediscovering the target model expression.
 Curiously, it performs best when the context is excluded and worst when the data is excluded.
-This suggests that the context we provided was counterproductive for the search, and purely reasoning about the data led greater success.
-We also note that GPT-4 finds the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$ when all operators are provided, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
-Kepler: We find strong evidence that the prompt has triggered GPT-4’s memorization of Kepler’s Law: the scratchpad reveals GPT-4 associates the variable names in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification.
-Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in GPT-4’s training data than the other relationships.
-The hard search for Kepler’s Law with all tools on has led GPT-4 to explore more complex expressions since the target model complexity is only 5 with MSE of 46.6886.
-GPT-4 also finds the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$.
-However, SciPy optimizes and fits $c_{2}$ with to a floating point power ($\approx 1.5$).
-We did not consider this to be a rediscovery of the target model as such expressions are not dimensionally consistent.
+This suggests that our context was counterproductive for the search, and purely reasoning about the data led to greater success, with GPT-4o leading the performance (Figure 8).
+Scratchpad shows that GPT-4o refers to Bode in a few experiments, specifically in relation to the ’semi-major axis’ variable identity provided in the context, which may have been overly specific.
+However, it never suggested the target model expression in those instances.
+We also notice GPT-4o rediscovering the target model expression in the first few iterations from just the data patterns (Figure 9).
+However, both GPT-4 and GPT-4o find the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
+Figure 8: GPT-4 and GPT-4o results on Bode’s data.
+For the “Hard Search” in the GPT-4 model, all the trends except ‘Without Context’ overlap at a score of 0.
+Transparency and different marker styles are used to differentiate these trends.
+Figure 9: GPT-4o reasoning on Bode’s data. The scratchpad snippet is from the first iteration in Run 4 with all tools on for hard search.
+Kepler: We find strong evidence that the prompt has triggered LLM’s memorization of Kepler’s Law: the scratchpad reveals that both GPT-4 and GPT-4o models associate the variable names and data in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification.
+Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in OpenAI’s training data than the other relationships.
+However, scratchpad also shows that although GPT-4o infers Kepler’s relationship from data and/or variable identity, it does not always suggest the target model expression but reasons with alternative expressions based on dataset patterns and feedback from prior suggestions.
+Both the models also often find the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$.
+However, SciPy optimizes and fits $c_{2}$ with a floating point power ($\approx 1.5$).
+We did not consider this a rediscovery of the target model as such expressions are not dimensionally consistent.
+Figure 10: GPT-4 and GPT-4o results on Kepler’s Third Law data.
+For the “Hard Search” in the GPT-4o model, the trends for ‘Without Data’ and ‘With All Tools’ overlap at a score of 0.
+Transparency and different marker styles are used to differentiate these trends.
 Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round.
-In easier searches, GPT-4 consistently found it within 15 iterations; inclusion of context and data seemed to improve performance.
-However, the hard search was much more challenging, with Langmuir only discovered once, in the case with all tools on.
+In easier searches, both GPT-4 and GPT-4o consistently found it within 15 iterations; inclusion of context and data seemed to improve performance.
+However, the hard search was much more challenging, with Langmuir discovered a few times, mainly with all tools on.
+An interesting observation about the GPT-4o model is that it references Langmuir and other adsorption models (Freundlich or BET), particularly for experiments with easy search and without data, where the model focuses more on the adsorption context.
+However, similar to other benchmark datasets, it does not always propose the target model expression; instead, it methodically analyzes the information it receives, including the context and feedback.
+Figure 11: GPT-4 and GPT-4o results on Langmuir’s adsorption data.
+For the “Hard Search” in the GPT-4 model, all the trends except ‘With All Tools’ overlap at a score of 0.
+Transparency and different marker styles are used to differentiate these trends.
 Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions.
 In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 12).
-However extrapolation shows this model is not theoretically correct (Figure 13 in SI).
-This dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92].
-Because this expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search.
+However, extrapolation shows this model is not theoretically correct (Figure 19 in SI).
+This may also explain why we see a decreasing score trend, even after identifying the target expression, compared to other datasets (Figure 13).
+We analyzed the LLM-generated expressions from both dual-site and Langmuir adsorption datasets by checking whether they satisfy thermodynamic constraints by running through a computer algebra system built in our previous work [28].
+We consistently found one more accurate expression at the expense of an added complexity for Henry’s Law term, which models adsorption at low pressures well.
+Figure 12: Pareto fronts for dual-site Langmuir dataset. The black line represents the best total front from the five independent runs. The target model is labeled as a blue star.
+Dual-site Langmuir dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92].
+Because the target model expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search in GPT-4 model.
 We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop).
-We found that when using basic operators and context, GPT-4 was getting close to finding the target model (Figure 8) upon running 15 iterations with this feedback loop.
+We found that when using basic operators (easy search) and all tools on, GPT-4 was getting close to finding the target model (Figure 12) upon running 15 iterations with this feedback loop.
 To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed.
 Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR.
-But when targeting models like dual-site Langmuir, these instructions may have been a liability, and we needed to adjust the prompt and feedback sent to GPT-4 to allow for the exploration of more complex expressions.
-Figure 8: Pareto fronts for dual-site Langmuir dataset. The black line represents the best total front from the five independent runs. The target model is labeled as a blue star.
-Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to GPT-4 exceeded the token limit.
+Figure 13: GPT-4 and GPT-4o on dual-site Langmuir’s adsorption data.
+All the trends in GPT-4 model overlap at a score of 0.
+Transparency and different marker styles are used to differentiate these trends.
+However, the same prompts were more successful with GPT-4o model, which reasoned better with data and context.
+For experiments without data, GPT-4o relied on the context of general adsorption behavior.
+It initiated with expressions based on Langmuir, Freundlich, and occasionally Temkin and BET theories.
+Interestingly, we observed the highest score on this dataset when no context was provided, and the target model was rediscovered through dataset patterns and iteration feedback.
+Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in the datasets related to saturation behavior and leading to modifications of the Langmuir expression to derive the dual-site behavior.
+Figure 14: Instances of GPT-4o reasoning on dual-site Langmuir dataset.
+The top snippet of the scratchpad is from Run 1, and the bottom is from Run 2.
+Both are for runs with all tools on and iteration 4.
+Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to LLMs exceeded the token limit.
 Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations.
-So, we developed a cost-saving scheme: only send a portion of the data in the prompt to GPT-4, while fitting and evaluating the generated expressions using the whole dataset (Figure 14 in SI).
-Since GPT-4 generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging.
+So, we developed a cost-saving scheme: only send a portion of the data in the prompt to LLM, while fitting and evaluating the generated expressions using the whole dataset (Figure 20 in SI).
+Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging.
 We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions.
 So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error.
-Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to GPT-4 for feedback.
+Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to LLMs for feedback.
 To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
 Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6].
-We modified the basic prompt to encourage GPT-4 to explore longer expressions and also tested the effect of “prodding" GPT-4 by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better.
-We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) random data points and the second (S2) with another random 36 (20%) data points.
+We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better.
+We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) data points and the second (S2) with another 36 (20%) data points.
 Overall, feeding in more data generated better-fitted and more complex expressions.
-Figure 9 shows the top expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^).
-Figure 9: Models for Nikuradse dataset from GPT-4.
-Here P refers to prompt versions, and S refers to dataset.
+Figure 15 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^).
+We also list the complexities and MAEs for these expressions in Table 1 in SI.
+Figure 15: Models for Nikuradse dataset.
+Here, P refers to prompt versions and S refers to dataset.
+The expressions generated by the GPT-4o model are marked with an ‘o’ at the end.
 We notice unphysical behavior with fewer data in prompt versions 1 and 2 (Fig.
 15 A-B) where GPT-4 is instructed to explore long expressions phrased in two different ways.
-However it generalizes better with fewer data for prompt version 3 where we challenge GPT-4 by providing information about the MAE and complexity of BMS model.
-We found an optimal expression from GPT-4 with a complexity of 41 and an MAE of 0.01086 (Figure 9 C).
-The MAE is approximately three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS).
+However, GPT-4o model only shows unphysical behavior in prompt version 1 with fewer data.
+Both the models generalize better with fewer data in prompt version 3 where we challenge by providing information about the MAE and complexity of the BMS model.
+We found optimal expressions from GPT-4 and GPT-4o with a complexity of 41 and 27 and an MAE of 0.01086 and 0.00924, respectively (Figure 15 C).
+GPT-4o performs qualitatively better, generating a more accurate expression at a significantly lower complexity than GPT-4.
+The MAE from both models is at least two to three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS).
 BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392.
-However unlike GPT-4, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps to identify this expression.
-Our GPT-4 model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions.
-We additionally assessed BMS on its default move probabilities and using 40 temperatures for parallel tempering as mentioned in the paper, and ran it for 153 MC steps with Nikuradse data -- it generated a constant function, and upon running it longer (1000 MC steps), it suggests an expression with MAE 0.13436 and complexity 25.
-Thus, we can see that GPT-4’s incremental suggestions for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS.
+However unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression.
+Our model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions.
+The optimal expression chosen is the best from 5 runs (1/153*5 or 1/765).
+The incremental suggestions from LLMs for new expressions (at least in the initial stages of the search) are of much higher quality than those by BMS.
 However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost.
-Using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
+For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
 BMS samples expressions from a probability distribution.
@@
 EFS is known to be highly efficient, delivering expressions in seconds.
-4 Discussion and Conclusion
+4 Discussion and Conclusions
 SR programs that optimize for speed aim to generate expressions quickly.
 In contrast, our proposed method emphasizes informed optimization, leveraging contextual information more effectively.
-The clearest cases of “leveraging the context” occurred when GPT-4’s first guess included the target model among three expressions.
+The clearest cases of “leveraging the context” occurred when the first guess included the target model among three expressions.
 But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions.
@@
 Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens.
+GPT-4o is an upgraded version of GPT-4 that offers improved performance, being both faster and more cost-effective [93].
+Consistent with observations made by others who have tested it on various tasks [94, 95], we also noted improvements in the quality of generated expressions and the effectiveness of scratchpad analysis through reasoning for symbolic regression.
 In general, we found natural language to be a rather clumsy interface for controlling expression length.
@@
 Classical approaches that incorporate expression length into measures of fitness or score are certainly more precise for controlling length, even if expression length is an imperfect measure of parsimony and meaningfulness in SR [96].
-We ran separate, focused tests to evaluate the effect of prompting on expression length, and GPT-4 did not obey instructions that requested, for example, "expressions of length 17." Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length.
-The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting complexity.
-Testing our approach on equation rediscovery using GPT-4 invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources.
-Indeed, the data are publicly available, as well, though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets.
-We found strong evidence of this because of our scratchpad implementation, which revealed when it was thinking of Kepler’s law before seeing the data.
+We ran separate, focused tests to evaluate the effect of prompting on expression length with the GPT-4 model.
+However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length.
+The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting operators and variables to calculate complexity.
+Testing our approach on equation rediscovery using language models invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources.
+Indeed, the data are publicly available, as well (appearing, for example, as tables in the .pdf documents of the historical papers), though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets.
+We found strong evidence of this when the scratchpad revealed the LLM naming Kepler’s law before seeing the data.
 Nonetheless, we found evaluating the model outputs to be informative.
 Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include.
-This context may include experimental details, instrument specifications, and literature, and even a “memorized” explanation of the data by an LLM or retrieved from context in a retrieval-augmented generation scheme [94] would be valuable.
-A true blind test would be to rediscover a novel scientific law using an LLM with a knowledge cutoff predating the seminal publication.
+Our experiments included $<$100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [97].
+The recent study by Shojaee and coworkers [37] demonstrated strong performance in generating scientific expressions using LLMs, using an approach that generated more structured outputs than our approach.
+Rather than rediscovering scientific laws as standalone expressions, they tasked the LLMs to “autocomplete” a structured Python function within the context of an optimization problem.
+They also designed new benchmark scenarios and datasets to highlight the strengths of this approach.
+This approach effectively leverages the strengths of language models in reasoning about well-defined problem statements.
+However, the extent to which its success can be attributed to prior exposure to similar questions during training remains unclear without a scratchpad to gain insights into its reasoning steps.
+A true blind test with these language models would be to rediscover an unpublished or recently-published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
 There are two ways to implement in-context learning in LLMs through prompts.
@@
 Data Availability
-The code, prompts, and data supporting the findings of this work are available at https://github.com/ATOMSLab/pySR_adsorption.
+The code, prompts, and data supporting the findings of this work are available at https://github.com/ATOMSLab/LLMsforSR.
 5 Supporting Information
 Figure 16: Illustration of GPT-3.5 attempting to perform SR.
-Figure 11: Context provided to GPT-4 for all the datasets
-Figure 12: Datasets explored using GPT-4 for SR
+Figure 17: Context provided to LLMs for all the datasets
+Figure 18: Datasets explored using LLMs for SR
 Figure 19: GPT-4-generated model does not follow theory on extrapolation. It gives a negative value for adsorption loading at lower pressures.
@@
 The red and yellow points represent the data sent to GPT-4 (selected randomly from the original dataset), while the grey ones show the original dataset sent to SciPy for optimization.
+Expression Mean Absolute Error (MAE) Complexity
+P1S1 $0.02270419$ $13$
+P1S2 $0.00978477$ $29$
+P2S1 $0.00897093$ $69$
+P2S2 $0.00931620$ $49$
+P3S1 $0.01086397$ $41$
+P3S2 $0.00992712$ $49$
+P1S1o $0.02007803$ $19$
+P1S2o $0.02141686$ $17$
+P2S1o $0.00954461$ $27$
+P2S2o $0.01186963$ $27$
+P3S1o $0.00923655$ $27$
+P3S2o $0.01144178$ $19$
+Table 1: Expressions of LLM models for Nikuradse dataset with complexity and MAE.
+Here, P refers to prompt versions and S refers to dataset.
+The expressions generated by the GPT4o model are marked with an ‘o’ at the end.

### v3

@@
 Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained.
-Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets.
+Here, we explore the potential of LLMs to perform symbolic regression – a machine-learning method for finding simple and accurate equations from datasets.
 We prompt GPT-4 and GPT-4o models to suggest expressions from data, which are then optimized and evaluated using external Python tools.
-These results are fed back to the models, which proposes improved expressions while optimizing for complexity and loss.
+These results are fed back to the LLMs, which propose improved expressions while optimizing for complexity and loss.
 Using chain-of-thought prompting, we instruct the models to analyze data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions.
-We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation.
-GPT-4 and GPT-4o successfully rediscovered all five equations and generally performed better when prompted to use a scratchpad and consider scientific context.
-GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse datasets.
+We evaluated the workflow in rediscovery of Langmuir and dual-site Langmuir’s model for adsorption, along with Nikuradse’s dataset on flow in rough pipes, which does not have a known target model equation.
+Both the GPT-4 and GPT-4o models successfully rediscovered equations, with better performance when using a scratchpad and considering scientific context.
+GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse dataset.
 We demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data.
-We also observe how theory can sometimes offset noisy data and, in other cases, data can make up for poor context.
+We also applied symbolic mathematical constraints based on the background knowledge of data via prompts and found that LLMs generate meaningful equations more frequently.
 Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
@@
 For scientific data, we often aim to find expressions with few adjustable parameters explaining the data while ensuring that they align with theory.
-Symbolic regression is a machine learning technique that approaches equation-based scientific discovery – given a dataset, it searches through some “space of possible equations” and identifies those that balance accuracy and simplicity.
+Symbolic regression is a machine learning technique that approaches equation-based scientific discovery.
+Given a dataset, it searches through some “space of possible equations" and identifies those that balance accuracy and simplicity.
 It is different from conventional regression methods, as symbolic regression infers the model structure from data rather than having a predetermined model structure.
-Mathematically, symbolic regression is formulated as some form of optimization, not just of the constants in an equation, but as a search through “equation space" for optimal expressions.
-In this way, symbolic regression is a form of machine learning – as data is received, an internal model is updated to match the data; when the model fits the data well and can make predictions about unseen data, the algorithm is said to have “learned" the underlying patterns in the data.
-In contrast to popular machine learning algorithms like neural networks, symbolic regression does not just only fit the constants in an equation but also finds functional forms that match the data.
+Mathematically, symbolic regression is formulated as some form of optimization, not just of the constants in an equation, but as a search through “equation space” for optimal expressions.
+In this way, symbolic regression is a form of machine learning – as data is received, an internal model is updated to match the data.
+When the model fits the data well and can make predictions about unseen data, the algorithm is said to have “learned” the underlying patterns in the data.
+In contrast to popular machine learning algorithms like neural networks, symbolic regression not only fits the constants in an equation but also finds functional forms that match the data.
 Symbolic regression methods mainly use genetic algorithms [59, 54, 86, 66] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached.
@@
 Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [36, 70, 68, 93, 40, 16, 56, 30, 55, 42, 92, 53, 72, 23, 33], including using large language models integrated with genetic algorithms [67, 61, 75, 11, 41, 43, 65, 73, 89, 38].
-As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
+As scientific data is strongly tied to theory, encoding it in the program narrows the vast search space and can make the programs more effective.
 This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery.
@@
 Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [75, 43], the context employed is limited to a list of previously-obtained expressions.
-We propose to expand the context to include data, as well natural language descriptions of the scientific context of the problem.
+We propose to expand the context to include data, as well as natural language descriptions of the scientific context of the problem.
 To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [6, 18, 98].
@@
 Symbolic regression requires equation generation and precise fitting of numerical constants.
-Yang and coworkers [68] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without showing the LLM the analytical form.
-They find that “LLM can often calculate the solution directly from the analytic form," however we find LLMs to be unreliable and inefficient for such tasks.
+Yang and coworkers [101] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without revealing the LLM the analytical form.
+They find that ‘LLM can often calculate the solution directly from the analytic form,’ however, we find LLMs to be unreliable and inefficient for such tasks [34].
 Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [41], iteratively refining prompts for more accurate output.
@@
 This approach aligns closely with two recent works [73, 89] in the literature where LLMs were used to generate expressions and optimizations were performed via iterations.
-Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using both data and context in natural language as model input, and implementing a scratchpad to record raw model output as it solves the problem.
+Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using data, context, and theoretical constraints in natural language as model input and implementing a scratchpad to record raw model output, as it solves the problem.
 This promotes informal reasoning about the data and context, and helps reveal “test set leakage.”
 2 Methods
-Preliminary tests
+2.1 Preliminary tests
 In our initial tests, we asked LLMs to generate mathematical expressions from scientific data.
-We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1).
+We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures.
 In all cases, the models produced expressions while “hallucinating” arbitrary coefficients.
-We revised the prompts to ask the LLMs to “show all steps" to gain insight into how the model selects these values [65].
-In response, the output either provided Python code for optimization or, at times, mathematical steps for optimization.
-Figure 1: Illustration of GPT-4 attempting symbolic regression.
-GPT-4 predicts expressions with optimized coefficients when passed a dataset for nitrogen adsorption on mica [69].
-The Python code snippet from Prompt 2 output has been truncated to keep the figure concise.
-Note that the actual parameter values produced by running the code differ from what GPT-4 generates.
-Figure 16 in SI shows the output from GPT-3.5 turbo which produces similar results.
-Although the generated code accurately performed curve-fitting, the LLM hallucinated incorrect coefficients.
-Nonetheless, the generated remarks about the data patterns hinted that they may have the potential to generate accurate functional forms that can be optimized outside the LLM.
-System design
-Therefore, we designed a workflow (Figure 2) where we task GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM.
-A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE).
-These results are stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions.
+We revised the prompts to ask the LLMs to “show all steps” to gain insight into how the model selects these values [98].
+In response, the output either provided Python code for optimization or at times, mathematical steps for optimization, but with the wrong output.
+Figure 1: GPT-4’s hallucinated curve fitting on adsorption data.
+The plot shows how GPT-4 predicts expressions with coefficients when passed a dataset for nitrogen adsorption on mica [60].
+The Python code that it generates is correct; however, the actual optimized parameter values produced by running the code differ from what GPT-4 outputs.
+Figure 19 and 20 in SI show the raw output from GPT-3.5 turbo and GPT-4.
+As illustrated in Figure 1, although the LLM-generated code runs and accurately performs curve-fitting, the model still hallucinates incorrect coefficients as its final output.
+Nonetheless, the generated remarks about the data patterns suggested that they may have the potential to generate accurate functional forms that can be optimized outside the LLMs.
+2.2 System design
+Therefore, we designed a workflow (Figure 2) where we task models GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM.
+A Python class takes in expressions, optimizes them, and then calculates their complexity and mean squared error (MSE).
+The results are then stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions.
 We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4 and GPT-4o, more frequently generating expressions that didn’t parse.
+Although LLMs could have been used to generate the curve-fitting code, we ensure reliability and prevent reward hacking by writing and validating the code ourselves.
+We define complexity as the number of operators and variables/constants in a given expression, or in other words, the total nodes in an expression tree (see Figure 3).
+A mathematical expression can be represented as a rooted tree structure where internal nodes correspond to operators, which can be binary (e.g., ++, *, //, or ˆ) or unary (like sin or log).
+The leaves, or terminal nodes, represent either variables (e.g., $x_{1}$, $x_{2}$) or constant values (e.g., $c_{1}$, $c_{2}$).
 Figure 2: Workflow for using LLMs for SR.
-First, the dataset is sent to LLMs (with or without context) which is instructed to suggest expressions without optimizing parameters.
+First, the dataset is sent to LLMs (with or without context), which are instructed to suggest expressions without optimizing parameters.
 The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [77] method with basin-hopping as the numerical optimizer [97]).
-Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions.
-The top performing expressions are sent as feedback to LLMs, which is asked to suggest better expressions optimizing for both complexity and loss.
+Results for each expression are stored in a Python dictionary and added to a list of results from prior expressions.
+The top-performing expressions are sent as feedback to LLMs, which are asked to suggest better expressions that optimize for both complexity and loss.
 The feedback loop is run for a set number of iterations.
+Figure 3: Langmuir’s adsorption model illustrated in expression tree.
+The equation ‘complexity’ is computed from the equation string: each binary operator (+, -, *, /, **) is counted and weighted by 2, each occurrence of a unary function name (e.g.
+log, exp, sqrt, sin, etc.) is counted with weight 1, and a base cost of 1 is added so that even a constant expression scores at least 1.
 We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 and GPT-4o to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format.
@@
 We implemented strategies like incorporating examples and a scratchpad and subsequently refined the prompts.
-Prompt Engineering
+2.3 Prompt engineering
 We prototyped our system by testing its ability to rediscover the Langmuir adsorption isotherm ($q=c_{1}*p/(c_{2}+p)$) from experimental data [60].
 This enabled us to quickly identify major structural improvements to the workflow; we further tailored the prompts while testing on more difficult problems.
-Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors.
+Removing bias: We aimed to make the workflow run smoothly without any human intervention, and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors.
 A simple way to illustrate the expected outcome was by providing examples in a few-shot context [14, 91].
@@
 While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
-Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax.
+Figure 4: Bias from prompt examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax.
 The revised prompts (Prompt 2 and 3) eliminate this bias from the equation prediction process.
@@
 To address this, we made a minor adjustment by removing the term "LaTeX" and allowing the model to generate expressions as Python strings instead.
-This small change was was sufficient to remove bias and was also less prone to syntax errors.
+This small change was sufficient to remove bias and was also less prone to syntax errors.
 We nervetheless processed the generated Python expressions through our formatting function, and it worked seamlessly.
@@
 After implementing this technique, the model immediately generated higher-quality expressions (Figure 5).
-(a) Effect of scratchpad on Langmuir Dataset
-(b) Effect of scratchpad on Kepler Law Dataset
-Figure 4: Illustration of scratchpad approach.
-We observe substantial, qualitative improvements in the predicted expressions after implementing the scratchpad technique (Prompt 2).
-The suggested expressions for (a) Langmuir’s and (b) Kepler’s Law dataset include operators (/// and sqrt, respectively) present in the target models ($y=\frac{c_{1}*x}{c_{2}+x}$ and $y^{2}=c_{1}x^{\frac{3}{2}}$, respectively)
+Figure 5: Effect of scratchpad on Langmuir Dataset.
+We observe substantial and qualitative improvements in the predicted expressions after implementing the scratchpad technique (Prompt 2).
+The suggested expressions for Langmuir’s dataset include operators (//) present in the target model ($y=\frac{c_{1}*x}{c_{2}+x})$.
 Preventing redundant expressions: The LLMs often generated expressions like $x+c_{1}$ and $x-c_{1}$, implying they are different.
@@
 SR programs often successfully generate expressions that fit the data well and are simple, yet they may not adhere to scientific principles or be otherwise “meaningful.” Yet scientists often have valuable insights into their domain that extend beyond these constraints, and they may not always know which specific expressions will best capture the nuances of their dataset or if entirely new expressions might be more effective.
-By incorporating scientific context, we aim to align the equation search to be consistent with scientific theories.
+By incorporating scientific context, we aim to align and constrain the equation search to be consistent with scientific theories.
 Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g.
 by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [85, 3, 29, 42, 68, 92, 12].
-Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 17 in SI lists the context provided for each of the datasets.
+Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations” and “consider scientific context.” Figure 21 in SI lists the context provided for each of the datasets.
 Figure 7: Illustration of GPT-4 following restrictions from prompt.
 3 Results
-We evaluated our workflow using experimental datasets associated with meaningful scientific context.
-SR benchmarks often use synthetic data; we designed our tests around a benchmark [88] specifically curated for evaluating SR algorithms for scientific data.
-From here, we selected three experimental datasets from astronomical observations (Bode’s Law, Hubble’s Law, and Kepler’s Law).
-To these, we added two experimental chemistry datasets: Langmuir and Dual-site Langmuir adsorption isotherm models [9].
-Figure 18 in SI illustrates all the datasets with the target expression for each.
-We also tested our workflow using a dataset on friction losses in pipe flow [89].
-This phenomena doesn’t have an established target model, and has been the subject of prior study by other SR programs [6].
-We included the whole dataset for each problem in the prompt to LLMs at the start and in each iteration.
-Because the Hubble’s dataset has an unreasonable number of digits following each entry, we rounded to 3 decimal places when sending data to LLMs but used the original dataset when optimizing constants in SciPy.
-This reduced the tokens (and cost) of running the models.
-We pass the entire dataset SR is different than other machine learning models, as we feed the entire dataset for training where the test is the output expression [88].
-We conducted eight sets of experiments for each model (GPT-4 and GPT-4o) on six datasets, running each set five times to evaluate its robustness.
-In four of the experiments, we used basic binary operators ($+$, --, *, $\div$), incorporating ‘sqrt’ for Kepler’s data and ‘^’ and ‘exp’ for Bode’s data.
-Empirical relations often involve field-specific operators, so the equation search should account for this.
-We refer to this set of tests as an “easy search" because the space of possible equations is constrained to that generated by the operators in our dataset.
-In addition, we conducted four further tests adding common unary operators (sqrt, log, exp, square, cube) alongside the basic ones to evaluate a more difficult search.
-All the datasets were run in four settings: without context, without data, without scratchpad and finally with all tools on.
-In the following text, we discuss each dataset individually and analyze their performance.
-The score at each iteration is a total of x/5 independent runs where target model expressions were found.
-Temperature is a hyperparameter used in stochastic models like LLMs to regulate the randomness of the model output [90, 91].
+We focused on applying LLM-based SR to experimental datasets, exploring its potential to uncover interpretable scientific relationships rather than benchmarking on synthetic datasets.
+Specifically, we investigated two well-known chemistry datasets: Langmuir and dual-site Langmuir adsorption isotherm models [60].
+These two datasets were also examined in our previous work [23, 33], which enables us to directly compare this LLM-based SR approach with traditional SR methods.
+We further tested our workflow on a challenging dataset of friction losses in pipe flow [78], which does not have an established target model and has been the subject of prior studies by other SR programs [40, 13].
+We used Nikuradse data from [40], which was sourced directly from the original reported literature [78], without any modifications.
+We also tried datasets corresponding to Hubble’s Law, Kepler’s Law, and Bode’s Law (from PySR’s empirical benchmark [25]).
+However, these were highly susceptible to different forms of leakage, and we therefore include those results only in the SI section 5.2 for interested readers.
+The scientific context was often sufficient for the LLMs to infer the correct law name and the target model expression in the first iterations.
+Additionally, the data, such as the orbital periods of the planets in the solar system, were recognizable to the LLMs even without the scientific context, as it appears in the Wikipedia article on Kepler’s Law [21].
+Figure 22 in SI illustrates all the explored datasets with their respective target model expressions.
+3.1 Experimental settings
+We conducted experiments using OpenAI’s GPT-4 and GPT-4o models in ‘Easy Search’ and ‘Hard Search’ settings.
+‘Easy Search’ includes only the basic binary operators (++, -, *, $\div$).
+In addition to this, the ‘Hard Search’ includes standard unary operators (sqrt, log, exp, square, cube) to evaluate a more challenging search space.
+Additionally, we compared the performance of LLMs on following mathematical constraints (derived from background theory of data) with our previously explored work using computer algebra systems [33].
+Our workflow has three main components:
+  • Data: sent as input in natural language.
+For instance, Langmuir’s data in the prompt looks like: “Your job is to find 3 expressions that approximately describe the dataset with dependent variable, y: [33, …] and independent variable, $x_{1}$: [34, …]”
+  • Context: is sent as a single-line text input about data and its variable identities; for instance, the context used for Langmuir’s data is: The data is about nitrogen adsorbing onto mica, where the independent variable (x1) is pressure, and the dependent variable (y) is loading.
+  • Scratchpad: frames the task as a reasoning problem in the context of data and free-form background information.
+For example, a scratchpad output from the LLMs for Langmuir’s data looks like: Looking at the dataset, we can see that ‘y decreases as $x_{1}$’ increases.
+Therefore, it seems to have an inverse relationship.
+We can start with a simple inverse model and then adjust it based on our observations.
+Temperature is a dimensionless hyperparameter used in stochastic models such as large language models (LLMs) to regulate the randomness of output generation [2, 1].
 It adjusts the probabilities of the predicted words in the softmax output layer of the model.
-Lowering temperature favors words with higher probability, so when the model randomly samples the next word from the probability distribution, it will be more likely to choose a more predictable response.
-We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the datasets.
-Hubble: Hubble’s Law is represented by the simple equation y = c1*x, which was often a first guess in the GPT-4 model when few operators are made available.
-However, including context plays a role; this is especially apparent in the “hard” search, where removing the scratchpad for reasoning or context inhibited performance (Figure 7).
-This dataset is particularly noisy, and including the noisy data as context actually inhibits the search as observed by consistent scores of 5/5 from both models when they were run without the data.
-Furthermore, the scratchpad in GPT-4o reveals that when the model first guesses Hubble’s Law, it then tends to reason and deduce a more complex, non-linear model in context of the noisy data.
-Figure 7: GPT-4 and GPT-4o results on Hubble’s data.
-In the GPT-4 model, all the trends except ‘Without Context’ overlap at score 5 for the “Easy Search” plot, and the trends ‘With All Tools’ and ‘Without Data’ overlap at score 4 for the “Hard Search”.
-In the GPT-4o model, the trends for ‘Without Data’ and ‘Without Context’ overlap at score 0 for the “Easy Search” plot, and all the trends except ‘Without Data’ overlap at score 0 for the “Hard Search”.
-Transparency and different marker styles are used to differentiate these trends.
-Bode: The results for Bode’s Law indicate that similar to other SR programs [88], LLMs encounter difficulty in accurately rediscovering the target model expression.
-Curiously, it performs best when the context is excluded and worst when the data is excluded.
-This suggests that our context was counterproductive for the search, and purely reasoning about the data led to greater success, with GPT-4o leading the performance (Figure 8).
-Scratchpad shows that GPT-4o refers to Bode in a few experiments, specifically in relation to the ’semi-major axis’ variable identity provided in the context, which may have been overly specific.
-However, it never suggested the target model expression in those instances.
-We also notice GPT-4o rediscovering the target model expression in the first few iterations from just the data patterns (Figure 9).
-However, both GPT-4 and GPT-4o find the target model as — $c_{1}\times\text{exp}(c_{2}\times x_{1})+c_{3}$, which is a symbolically-equivalent way of expressing Bode’s Law, though lacking the interpretability of the original form.
-Figure 8: GPT-4 and GPT-4o results on Bode’s data.
-For the “Hard Search” in the GPT-4 model, all the trends except ‘Without Context’ overlap at a score of 0.
-Transparency and different marker styles are used to differentiate these trends.
-Figure 9: GPT-4o reasoning on Bode’s data. The scratchpad snippet is from the first iteration in Run 4 with all tools on for hard search.
-Kepler: We find strong evidence that the prompt has triggered LLM’s memorization of Kepler’s Law: the scratchpad reveals that both GPT-4 and GPT-4o models associate the variable names and data in the context with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification.
-Perhaps because this relationship is routinely taught in high school and college physics courses, and thus likely to be more represented in OpenAI’s training data than the other relationships.
-However, scratchpad also shows that although GPT-4o infers Kepler’s relationship from data and/or variable identity, it does not always suggest the target model expression but reasons with alternative expressions based on dataset patterns and feedback from prior suggestions.
-Both the models also often find the target model in the form $c_{1}x_{1}^{c_{2}}$, with $c_{2}$ left as a constant for optimization instead of $\frac{3}{2}$.
-However, SciPy optimizes and fits $c_{2}$ with a floating point power ($\approx 1.5$).
-We did not consider this a rediscovery of the target model as such expressions are not dimensionally consistent.
-Figure 10: GPT-4 and GPT-4o results on Kepler’s Third Law data.
-For the “Hard Search” in the GPT-4o model, the trends for ‘Without Data’ and ‘With All Tools’ overlap at a score of 0.
-Transparency and different marker styles are used to differentiate these trends.
-Langmuir: Langmuir’s model is more obscure, and was almost never guessed in the first round.
-In easier searches, both GPT-4 and GPT-4o consistently found it within 15 iterations; inclusion of context and data seemed to improve performance.
-However, the hard search was much more challenging, with Langmuir discovered a few times, mainly with all tools on.
-An interesting observation about the GPT-4o model is that it references Langmuir and other adsorption models (Freundlich or BET), particularly for experiments with easy search and without data, where the model focuses more on the adsorption context.
-However, similar to other benchmark datasets, it does not always propose the target model expression; instead, it methodically analyzes the information it receives, including the context and feedback.
-Figure 11: GPT-4 and GPT-4o results on Langmuir’s adsorption data.
-For the “Hard Search” in the GPT-4 model, all the trends except ‘With All Tools’ overlap at a score of 0.
-Transparency and different marker styles are used to differentiate these trends.
-Dual-site Langmuir: The dual-site Langmuir model is particularly challenging for SR, because the target model does not significantly fit the data much better than many shorter expressions.
-In fact, GPT-4 found one expression that fits the data more accurately at complexity 11 (Figure 12).
-However, extrapolation shows this model is not theoretically correct (Figure 19 in SI).
-This may also explain why we see a decreasing score trend, even after identifying the target expression, compared to other datasets (Figure 13).
-We analyzed the LLM-generated expressions from both dual-site and Langmuir adsorption datasets by checking whether they satisfy thermodynamic constraints by running through a computer algebra system built in our previous work [28].
-We consistently found one more accurate expression at the expense of an added complexity for Henry’s Law term, which models adsorption at low pressures well.
-Figure 12: Pareto fronts for dual-site Langmuir dataset. The black line represents the best total front from the five independent runs. The target model is labeled as a blue star.
-Dual-site Langmuir dataset was previously explored in literature [9, 28] with three SR algorithms: Bayesian-based SR (BMS) [6], genetic programming-based SR (PySR) [88], and mixed-integer non-linear programming-based SR [92].
-Because the target model expression is longer, we extended the run to 50 iterations, and dual-site Langmuir was still not found for any of the experimental settings with easy search in GPT-4 model.
-We modified the feedback loop passed into GPT-4 for this dataset to send more accurate and longer expressions in lieu of passing the entire Pareto front (keeping the top five expressions based on MSE in the loop).
-We found that when using basic operators (easy search) and all tools on, GPT-4 was getting close to finding the target model (Figure 12) upon running 15 iterations with this feedback loop.
-To investigate this further, we ran with basic operators (for runs with and without context) for 50 iterations and found the target model in 1/5 runs for both tests, indicating that more iterations were needed.
-Our prompts instruct GPT-4 to generate shorter expressions, so it outputs simpler expressions with reasonable MSEs; parsimony is a common goal in SR.
-Figure 13: GPT-4 and GPT-4o on dual-site Langmuir’s adsorption data.
-All the trends in GPT-4 model overlap at a score of 0.
-Transparency and different marker styles are used to differentiate these trends.
-However, the same prompts were more successful with GPT-4o model, which reasoned better with data and context.
-For experiments without data, GPT-4o relied on the context of general adsorption behavior.
+Lowering the temperature favors words with higher probability, so when the model randomly samples the next word from the probability distribution, it will be more likely to choose a more predictable response.
+We tested the Langmuir dataset with five different temperature settings (0, 0.3, 0.5, 0.7, 1) and found 0.7 to be performing the best, which we later used for the rest of the experiments.
+Henceforth, the results are divided into four parts –
+  • Section 3.2 presents the results from Langmuir’s adsorption data, using Pareto fronts and ablation studies on each experimental factor.
+  • Section 3.3 provides analogous results for the dual-site Langmuir dataset.
+  • Section 3.4 compares the quality of LLM-generated expressions with constraints added from background knowledge on both Langmuir and dual-site Langmuir datasets.
+  • Section 3.5 demonstrates the applicability of this novel method on a larger dataset through the Nikuradse dataset [78], which is a multivariate experimental dataset that correlates friction with pipe roughness and Reynolds’ number, comprising over 350 measurements.
+3.2 Langmuir’s adsorption model
+The Pareto fronts below illustrate the overall performance of the GPT-4 and GPT-4o models in both ‘Easy Search’ and ‘Hard Search’ settings.
+Each experiment was conducted five times, and the expressions from these five runs are represented as scatter points in five different colors.
+Each Pareto front is also shown in its respective color, with an overall Pareto front indicated by black dashed lines.
+The overall front is created by pooling the results from all five Pareto fronts to present a single representation of the best performance across all runs.
+No model is less complex or has less MSE than the one model in the Pareto frontier.
+The target model expression is marked with a blue star, and its corresponding complexity and mean squared error (MSE) are indicated by grey dashed lines across the axes.
+As shown in Figure 8, the target model expression for Langmuir was identified in all cases, except for the ‘Hard Search’ with GPT-4.
+Figure 8: Pareto fronts for Langmuir’s adsorption model. The target model is expressed as $\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}$ with an MSE of 0.12263 and complexity of 7 (blue star).
+We also performed ablation studies under four configurations: without context, without data, without scratchpad, and with all tools enabled (the latter corresponding to the Pareto Fronts shown in 8).
+Figure 9 analyzes the performance of both models at each iteration level for all experimental settings.
+Figure 9: Ablation studies on Langmuir’s adsorption data.
+The score at each iteration represents the number of runs (out of five) in which the models successfully recovered the target model expression.
+Langmuir’s model is obscure and was hard to discover in the first rounds when no context was provided.
+Without data, the models rely more on the adsorption context, which increases the likelihood of accurately inferring the target model expression early on.
+However, while the models reference Langmuir, they also refer to other popular adsorption models like Freundlich or BET, and do not always propose the target model expression without considering the dataset pattern (Figure 10).
+Figure 10: Instances of GPT-4o reasoning on Langmuir dataset.
+The snippet of the scratchpad is from an experiment in Easy Search and without data (run #1\#1, iteration #1\#1).
+The purple-highlighted text shows references to popular adsorption models, the yellow ones show general observations about context and data, and the blue-highlighted one shows the generated target model equation.
+3.3 Dual-site Langmuir’s adsorption model
+Figure 11: Pareto fronts for dual-site Langmuir’s adsorption model.
+The target model is expressed as $\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}+\frac{c_{1}\times x_{1}}{c_{1}+x_{1}}$ with an MSE of 0.00079844 and complexity of 15 (blue star).
+Figure 12: Ablation studies on dual-site Langmuir’s adsorption data.
+The score at each iteration represents the number of runs (out of five) in which the models successfully recovered the target model expression.
+Figure 13: Extrapolation reveals theoretical
+inconsistencies of models.
+Dual-site Langmuir’s model dataset is particularly challenging for SR because the target model does not significantly fit the data much better than many other shorter expressions.
+This was also observed in previous literature studies [23, 33] using three SR algorithms: Bayesian-based SR (BMS) [40], genetic programming-based SR (PySR) [25], and mixed-integer nonlinear programming-based SR [4].
+As seen in Figure 11, the models generate other expressions with similar accuracy and half the complexity that reach almost similar accuracy as the target model expressions, creating a plateau from complexity 7 onward.
+However, most of these shorter expressions are inconsistent with the theory.
+For instance, extrapolation of such an expression (with complexity 11) generates a negative value for adsorption loading at lower pressures (Figure 13).
+Our prompts instruct LLMs to generate simple expressions, so they output shorter equations with reasonable MSEs; parsimony is a common goal in SR.
+Since the target model expression is longer for this dataset, we extended the run to 50 iterations and modified the feedback loop to pass only the top five expressions ranked by MSE in order to send more accurate and longer expressions in the feedback.
+This led to a decline in scores over iterations when the model initially discovered the ground-truth expression but later discarded it, as newer expressions with lower MSE replaced it in the top five, as shown in the ablation studies in Figure 12.
+This was also observed in case of Langmuir’s data (Figure 9) with GPT-4o.
+However, this does not affect the overall outcome since the best expressions that balance both accuracy and complexity were also evaluated in the Pareto fronts, as shown in Figures 8 and 11.
+We did not perform experiments with ‘Hard Search’ for this dataset, as the ‘Easy Search’ case already presented significant challenges, especially for GPT-4 model.
+GPT-4o model was more successful and reasoned better with data, context, and scratchpad.
+For experiments without data, like in Langmuir’s case, the models relied on the context of general adsorption behavior.
 It initiated with expressions based on Langmuir, Freundlich, and occasionally Temkin and BET theories.
-Interestingly, we observed the highest score on this dataset when no context was provided, and the target model was rediscovered through dataset patterns and iteration feedback.
-Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in the datasets related to saturation behavior and leading to modifications of the Langmuir expression to derive the dual-site behavior.
+Figure 14 presents two snippets from the scratchpad, illustrating how the dual-site Langmuir model was rediscovered by analyzing patterns in datasets related to saturation behavior, leading to modifications of the Langmuir expression to derive the dual-site behavior.
 Figure 14: Instances of GPT-4o reasoning on dual-site Langmuir dataset.
-The top snippet of the scratchpad is from Run 1, and the bottom is from Run 2.
-Both are for runs with all tools on and iteration 4.
-Nikuradse: Finally, to test the scalability of the approach to larger datasets and to test a problem without a known target model, we evaluated the Nikuradse dataset, which is experimental data for turbulent friction in rough pipes conducted by Johann Nikuradse in the early 1930s [89] The Nikuradse dataset contains over 350 measurements; including the whole dataset in our prompts to LLMs exceeded the token limit.
-Even with long context windows, analyzing large datasets would be expensive, since each iteration is more costly, and because generating longer expressions requires more iterations.
-So, we developed a cost-saving scheme: only send a portion of the data in the prompt to LLM, while fitting and evaluating the generated expressions using the whole dataset (Figure 20 in SI).
-Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimizing was also more challenging.
+The top snippet of the scratchpad is from run 1 (iteration # 5), and the bottom is from run 2 (iteration # 4).
+Both are for experiments with all tools.
+The purple highlighted text shows the reference to Langmuir adsorption model, yellow ones show the general observations about context and data pattern, and blue highlighted one shows the generated target model equation.
+3.4 Constraining equation search with natural language prompts
+SR algorithms traditionally explore the equation search space randomly, which is inefficient for scientific datasets following background knowledge, often leading to inconsistent equations, as seen in the dual-site Langmuir case.
+In our previous work [33], we used a computer algebra system to limit the search space of two SR algorithms, PySR (genetic algorithm-based) and BMS (Markov chain Monte Carlo-based), by applying symbolic mathematical constraints.
+This approach allowed us to generate expressions that were meaningful and theoretically consistent.
+However, incorporating symbolic constraints is challenging in an existing SR package as it requires software development across multiple layers.
+In this work, we leverage LLMs to simplify this process, and we express these constraints in natural language in the prompts, which is far simpler and reduces human effort.
+The earlier study, with Fox et.
+al [33], also examined single-component adsorption datasets, two of which overlap with the datasets explored in this work – Langmuir’s and dual-site Langmuir’s adsorption models.
+The constraints are formulated as follows:
+$\lim_{p\to 0}f(p)=0$ (1)
+$\lim_{p\to 0}f^{\prime}(p)<\infty$ (2)
+$\forall p>0\qquad f^{\prime}(p)\geq 0$ (3)
+Here, Constraint 1 ensures that all molecules must desorb, and loading cannot be negative in the limit of zero pressure.
+Constraint 2 requires that the slope of the isotherm must be a positive finite constant in the limit of zero pressure.
+And, Constraint 3 requires that loading does not decrease with increasing pressure (the isotherm is monotonically non-decreasing) for all ($\forall$) positive values of pressure.
+We constructed the same constraints in our prompts:
+  1. 1.
+    Passes through origin: as $x1\rightarrow 0^{+}$, $y(x1)\rightarrow 0$.
+  2. 2.
+    Finite, positive initial slope: as $x1\rightarrow 0^{+}$, $0<dy/dx_{1}<\infty$
+  3. 3.
+    Monotonic non-decrease: for all $x1>0,dy/dx1\geq 0$
+We demonstrate the effect of adding constraints on the generated expressions in the GPT-4o model in comparison with PySR and BMS in Figure 15.
+(a) Effect of constraints on Langmuir dataset.
+(b) Effect of constraints on the dual-site Langmuir dataset.
+Figure 15: Comparison of generated expressions with constraints on and off conditions across GPT-4o, PySR, and BMS methods.
+The data for PySR and BMS is as reported in [33].
+The data for GPT-4o is from the ‘Easy Search’ experiments to align with the operators used in both PySR and BMS, as well as with all tools.
+For each run, the passing rate was calculated as the proportion of expressions (across all iterations within that run) satisfying each constraint.
+Error bars represent the standard error of the mean across the five runs.
+Notably, the largest improvements are observed with Constraint 1 across both datasets (see Figures 15 and 16).
+Despite its simplicity, the training data may not accurately reflect the underlying thermodynamic principles.
+An expression that fits the data well but does not approach zero loading during interpolation is unphysical and should not appear on the Pareto front as a top-performing one.
+The effect becomes more noticeable for the dual-site Langmuir case, and the benefit of constraining the search space becomes clearer, as the target model achieves accuracy comparable to shorter expressions that violate domain-specific constraints (see Figure 13).
+(a) Distribution of root mean square error (RMSE) for Langmuir expressions.
+(b) Distribution of root mean square error (RMSE) for dual-site Langmuir expressions.
+Figure 16: Distribution of generated expressions with constraints on and off conditions across the GPT-4o model.
+Green areas represent models that passed constraint C1, while orange areas represent models that failed across the five iterations.
+The dashed line indicates the RMSE of the target model, and the y-axis shows the percentage of total expressions in each category.
+3.5 Performance of LLM-SR on open scientific problems
+Finally, to test the scalability of the approach to larger and multivariate datasets, and to evaluate a problem without a known target model in literature, we examined the Nikuradse dataset, which is experimental data on turbulent friction in rough pipes by Johann Nikuradse in the early 1930s [78].
+The data correlates the friction factor ($\lambda$) with relative roughness (k/D) and Reynolds’ number (Re), and contains over 350 measurements.
+Including the entire dataset in our prompts to LLMs exceeded the token limit, and even with long context windows, analyzing large datasets would be expensive, as each iteration is more costly, and generating longer expressions requires more iterations.
+Therefore, we developed a cost-saving scheme: only send a portion of the data in the prompt to the LLM, while fitting and evaluating the generated expressions using the entire dataset (Figure 26).
+Since LLMs generated longer and more complex expressions for this dataset (seven or more fitted constants were common), numerical optimization was also more challenging.
 We found the optimized coefficients varied slightly due to stochasticity in the basin-hopping algorithm; this could lead to inaccurate sorting of the generated expressions.
-So, we optimized constants ten times for each expression, then selected that with the lowest mean absolute error.
+So, we optimized the constants ten times for each expression, then selected the one with the lowest mean absolute error (MAE).
+We evaluate performance on this dataset computing for MAE to align with the performance metrics used in the literature by other SR programs.
 Additionally, we stored the fitted parameters in the feedback loop to assess the expressions sent to LLMs for feedback.
 To manage the context window and encourage longer expressions, we used the modified feedback loop that proved modestly successful for the dual-site Langmuir dataset.
-Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs in the literature[6].
-We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing just the MSE achieved by a literature model (without leaking the model), and challenging it to do better.
+Though there is no definitive target model for the Nikuradse data, we compared other candidate model expressions from different SR programs discussed by Guimerà, et al [40].
+We modified the basic prompt to encourage LLMs to explore longer expressions and also tested the effect of “prodding" by sharing the MAE achieved by a literature model (without leaking the model), and challenging it to do better.
 We conducted six experiments on Nikuradse data, with three slightly different versions of the prompt (P1,P2, and P3) and two sets (S1 and S2) of data points — one (S1) with 36 (10%) data points and the second (S2) with another 36 (20%) data points.
 Overall, feeding in more data generated better-fitted and more complex expressions.
-Figure 15 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators ($+$, --, $\div$, $\times$ and ^).
-We also list the complexities and MAEs for these expressions in Table 1 in SI.
+Figure 17 shows the best expressions with the lowest MAE out of the six experiments that were explored with binary math operators (++, -, $\div$, $\times$ and ˆ).
+We also list the complexities and MAEs for these expressions in Table 1.
 Figure 17: Models for Nikuradse dataset.
-Here, P refers to prompt versions and S refers to dataset.
+Here, P refers to prompt versions and S refers to the dataset.
 The expressions generated by the GPT-4o model are marked with an ‘o’ at the end.
@@
 However, GPT-4o model only shows unphysical behavior in prompt version 1 with fewer data.
-Both the models generalize better with fewer data in prompt version 3 where we challenge by providing information about the MAE and complexity of the BMS model.
+Both the models generalize better with fewer data in prompt version 3, where we challenge by providing information about the MAE and complexity of the BMS model.
 We found optimal expressions from GPT-4 and GPT-4o with a complexity of 41 and 27 and an MAE of 0.01086 and 0.00924, respectively (Figure 17 C).
 GPT-4o performs qualitatively better, generating a more accurate expression at a significantly lower complexity than GPT-4.
-The MAE from both models is at least two to three times worse than the top-performing model identified by the Bayesian Machine Scientist (BMS).
-BMS uses Markov Chain Monte Carlo (MCMC)-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392.
-However unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression.
+The MAE from both models is at least two to three times worse than the top-performing model identified by the BMS.
+BMS uses MCMC-based SR and discovered a more accurate expression at complexity 37 with an MAE of 0.00392.
+However, unlike the language models, BMS evaluates thousands of expressions in more than 18000 Monte Carlo (MC) steps with parallel tempering to identify this expression.
 Our model uses a portion of the data to find the best expression from only a pool of (50*3 + 3) 153 expressions.
@@
 However, BMS is far more efficient in terms of compute, it generates more expressions at a far lower cost.
-For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for 5 runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
+For instance, using GPT-4 to generate 18000 trial expressions would be far too expensive (the cost of GPT-4 API calls for five runs of our workflow was about \$27 to obtain 153 expressions with the larger selected data points).
 BMS samples expressions from a probability distribution.
 Therefore, running for a long time is expected to be characterized by equilibration, in which the expressions converge to a region in the stationary distribution of expressions, after which no significant improvement would be observed apart from continuing to explore the region of most likely expressions.
-Genetic algorithms like Eureqa [3] do not converge to a stationary distribution and nor does our approach.
+Genetic algorithms like Eureqa [86] do not converge to a stationary distribution, nor does our approach.
 On the contrary, running a chatbot for a very long time leads to degradation of its answers as the context gets longer [63, 88], so, in principle, our method may exhibit similar degradation with very long runs.
 However, we expect this effect to be mitigated because we discard the majority of the context after each iteration, only passing the top and recent expressions.
-We also show a model from Eureqa in [6] which was run with default operator penalties and selected for the best expression from at least $10^{13}$ ones.
-The result from the EFS [85] model (best fit selected from 100 runs) in [6] is comparable to our model with an MAE of 0.00941.
-EFS is based on sparse regression that uses a genetic algorithm to generate basis functions automatically.
-EFS is known to be highly efficient, delivering expressions in seconds.
+Figure 18: Nikuradse models across the state-of-the-art symbolic regression programs.
+Here we compare our best suggested model from GPT-4o with other candidate models suggested from BMS, Eureqa, and EFS programs as reported in [40].
+All three methods’ featured models are at different points on the Pareto Front; GPT-4o generated few equations in the regime where Eureqa / BMS found optimal equations.
 4 Discussion and Conclusions
@@
 The clearest cases of “leveraging the context” occurred when the first guess included the target model among three expressions.
-But even when the search took longer, we found incorporation of the context, data, and scratchpad to be helpful for improving the quality of generated expressions.
+But even when the search took longer, we found that incorporating the context, data, and scratchpad helped improve the quality of generated expressions.
 However, including noisy data in the context sometimes undermined the search, as did including lower-quality scientific context.
-Nonetheless, this comes with great computational expense - especially since large datasets and long reasoning chains require so many tokens.
+Nonetheless, this comes with great computational expense, especially since large datasets and long reasoning chains require so many tokens.
 GPT-4o is an upgraded version of GPT-4 that offers improved performance, being both faster and more cost-effective [48].
 Consistent with observations made by others who have tested it on various tasks [76, 102], we also noted improvements in the quality of generated expressions and the effectiveness of scratchpad analysis through reasoning for symbolic regression.
-In general, we found natural language to be a rather clumsy interface for controlling expression length.
+In general, we found natural language to be a somewhat clumsy interface for controlling expression length.
 Different prompts and feedback mechanisms led to distributions of expressions with varying length.
@@
 We ran separate, focused tests to evaluate the effect of prompting on expression length with the GPT-4 model.
-However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure complexity of an expression accurately and to generate expressions of the target length.
+However, it did not obey instructions that requested, for example, “expressions of length 17.” Even with a scratchpad available, it failed to both measure the complexity of an expression accurately and to generate expressions of the target length.
 The LLM is better-suited for creative generation, while deterministic Python tools are more effective (and cheaper) at procedural tasks such as counting operators and variables to calculate complexity.
 Testing our approach on equation rediscovery using language models invariably involves a form of “test set leakage” since these expressions are on Wikipedia and countless additional Internet sources.
-Indeed, the data are publicly available, as well (appearing, for example, as tables in the .pdf documents of the historical papers), though we think it unlikely that LLMs trained on natural language would devote a sufficient fraction of their network to memorize these datasets.
-We found strong evidence of this when the scratchpad revealed the LLM naming Kepler’s law before seeing the data.
-Nonetheless, we found evaluating the model outputs to be informative.
-Moreover, we foresee a use case for scientists trying to solve a mystery about their data while having a great deal of context to potentially include.
-Our experiments included $<$100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [97].
+While the data are publicly accessible, for instance, in tables within historical research papers in PDF format, we think it is unlikely that LLMs trained primarily on natural language would devote a significant portion of their network to memorize these specific datasets.
+We found strong evidence when our prompts triggered LLM’s memorization of Kepler’s Law, where scratchpad revealed that both GPT-4 and GPT-4o models associate the variable names and data with Kepler’s Law, and it not only guesses the right answer in the first iteration, it names Kepler’s Law in its justification.
+Perhaps because this relationship is routinely taught in high school and college physics courses, and also the tabulated data is in Wikipedia, and thus likely to be more represented in OpenAI’s training data than others.
+In our analysis of adsorption datasets, we did not find such instances.
+The model proposed commonly referenced adsorption isotherms from the literature and recommended them based on patterns in the dataset; however, it never linked the raw data numbers to an adsorption model.
+Furthermore, the Nikuradse dataset provided no evidence of test set leakage.
+This is a use case we foresee when scientists are trying to solve a mystery about their data while having a great deal of context to include potentially.
+Our experiments included <<100 words of context, but this context could include experimental details, instrument specifications, and related literature, pushing the limits of the LLM’s context window, or going beyond the context window in a retrieval-augmented generation scheme [58].
 The recent study by Shojaee and coworkers [89] demonstrated strong performance in generating scientific expressions using LLMs, using an approach that generated more structured outputs than our approach.
@@
 However, the extent to which its success can be attributed to prior exposure to similar questions during training remains unclear without a scratchpad to gain insights into its reasoning steps.
-A true blind test with these language models would be to rediscover an unpublished or recently-published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
-There are two ways to implement in-context learning in LLMs through prompts.
-One is the Few-Shot prompting method, where we condition the model on a few specific examples related to the task that helps the model understand and perform the task more accurately.
-The other is the Zero-Shot prompting method, where the output relies solely on template-based prompts without specific task examples, allowing the model to infer how to handle the task from general instructions.
-Our approach does not provide examples of symbolic regression procedures within the prompts.
-Instead, we guide the model to engage in freeform chain-of-thought reasoning about the context as it prepares suggestions for new equations.
+A true blind test with these language models would be to rediscover an unpublished or recently published scientific law, using an LLM with a knowledge cutoff predating the seminal publication.
 Biasing the search space in a standard SR program can be challenging and requires advanced software and coding skills.
 Interdisciplinary work demands significant time and resources from researchers.
-Natural language interfaces in LLMs can help reduce some of these barriers by making program execution more accessible, even without expertise in software development.
+We demonstrate how natural language interfaces in LLMs can help reduce some of these barriers by making program execution more accessible, even without expertise in software development.
 Well-crafted prompts empower language models to perform diverse tasks, allowing them to adapt to different contexts and objectives.
 We can guide these models through prompts to generate responses that align with our specific needs, whether solving complex problems, developing creative content, or analyzing data.
+There are two ways to implement in-context learning in LLMs through prompts.
+One is the Few-Shot prompting method, where we condition the model on a few specific examples related to the task that help the model understand and perform the task more accurately.
+The other is the Zero-Shot prompting method (used here), where the output relies solely on template-based prompts without specific task examples, allowing the model to infer how to handle the task from general instructions.
+Our approach does not provide examples of symbolic regression procedures within the prompts.
+Instead, we guided the model to engage in free-form chain-of-thought reasoning about the context as it prepares suggestions for new equations.
 Acknowledgements
@@
 5 Supporting Information
+5.1 System Design
 Figure 19: Illustration of GPT-3.5 attempting to perform SR.
+Figure 20: Illustration of GPT-4 attempting symbolic regression.
+GPT-4 predicts expressions with optimized coefficients when passed a dataset for nitrogen adsorption on mica [60].
+The Python code snippet from Prompt 2 output has been truncated to keep the figure concise.
+Note that the actual parameter values produced by running the code differ from what GPT-4 generates.
 Figure 21: Context provided to LLMs for all the datasets
 Figure 22: Datasets explored using LLMs for SR
-Figure 19: GPT-4-generated model does not follow theory on extrapolation. It gives a negative value for adsorption loading at lower pressures.
+5.2 Additional results
+The following three datasets are from [25], carefully selected for evaluating SR algorithms for scientific data.
+The original data (Bode’s Law [10], Hubble’s Law [47], and Kepler’s Law [52]) were manually curated by [25] which we have used without modification.
+Figure 23: Pareto fronts for Bode’s Law.
+Figure 24: Pareto fronts for Hubble’s Law.
+Figure 25: Pareto fronts for Kepler’s Third Law.
 Figure 26: Nikuradse Dataset.
@@
 P3S2 $0.00992712$ $49$
-P1S1o $0.02007803$ $19$
+P1S1o $0.02007803$ 1919
 P1S2o $0.02141686$ $17$
@@
 P3S1o $0.00923655$ $27$
-P3S2o $0.01144178$ $19$
+P3S2o $0.01144178$ 1919
 Table 1: Expressions of LLM models for Nikuradse dataset with complexity and MAE.